# RILe: Reinforced Imitation Learning

### v3

@@ Abstract
Acquiring complex behaviors is essential for artificially intelligent agents, yet learning these behaviors in high-dimensional settings poses a significant challenge due to the vast search space. Traditional reinforcement learning (RL) requires extensive manual effort for reward function engineering. Inverse reinforcement learning (IRL) uncovers reward functions from expert demonstrations but relies on an iterative process that is often computationally expensive. Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons offer insufficient feedback for effective learning. We introduce RILe (Reinforced Imitation Learning), a framework that combines the strengths of imitation learning and inverse reinforcement learning to learn a dense reward function efficiently and achieve strong performance in high-dimensional tasks. RILe employs a novel trainer–student framework: the trainer learns an adaptive reward function, and the student uses this reward signal to imitate expert behaviors. By dynamically adjusting its guidance as the student evolves, the trainer provides nuanced feedback across different phases of learning. Our framework produces high-performing policies in high-dimensional tasks where direct imitation fails to replicate complex behaviors. We validate RILe in challenging robotic locomotion tasks, demonstrating that it significantly outperforms existing methods and achieves near-expert performance across multiple settings.
@@ 1 Introduction
Learning complex behaviors is critical for advancing artificially intelligent agents in fields such as robotics and strategic games. Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design, which is both time-consuming and error-prone.
@@
There are two ways to address the reward engineering problem. First, Inverse Reinforcement Learning (IRL) [1, 2] offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering. IRL proceeds iteratively: it first trains a policy (the learning agent’s decision-making mechanism) using the current reward function, observes how well the agent’s behavior aligns with the expert’s, and then refines the reward function to better guide the policy toward expert-like behaviors. Repeating this process eventually yields a reward function capable of providing nuanced feedback at different stages of learning. However, this iterative procedure is computationally expensive [3], especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
@@
Second, Imitation learning (IL) bypasses explicit reward design by directly comparing learned behaviors to expert demonstrations via a comparison mechanism. Traditional IL approaches such as Behavioral Cloning (BC) [4] match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks. To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL [5], introduce a discriminator that operates as a comparison mechanism and judges how expert-like the learned behaviors are. However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements. Instead, they rely on a comparison metric, often a distance measure or binary classification, that merely checks whether the agent’s behavior is (or is not) similar to the expert. Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize. Consequently, both traditional IL and AIL struggle in high-dimensional environments [6, 7], where the agent needs more granular and adaptive feedback than these mechanisms provide.
@@
Adversarial Inverse Reinforcement Learning (AIRL) [8] attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator. However, AIRL tightly couples the reward function to the discriminator’s output, causing it to inherit AIL’s limitations in high-dimensional settings where more fine-grained guidance is needed.
@@
Real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog. The teacher also refines how they teach as the student progresses. Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student’s learning progress more effectively. By contrast, existing approaches lack this continuous cooperative synergy. Adversarial Imitation Learning (AIL) does update a discriminator alongside the policy, but the discriminator’s sole role is to distinguish expert-like behavior from non-expert behavior. Consequently, the student attempts to fool this judge into classifying its behavior as expert-like, resulting in a competitive process rather than a cooperative trainer that dynamically shapes rewards based on suboptimal behaviors. Meanwhile, IRL methods only refine the reward after the policy converges, missing the opportunity for continuous co-evolution throughout training.
@@
To address these issues, inspiring by these insights, we propose Reinforced Imitation Learning (RILe). RIle combines the adaptive reward benefits of IRL with the computational efficiency of AIL (Fig. 1-(d)). RILe is a novel trainer-student system that establishes a positive-sum relationship between the trainer and the student. Specifically, RILe is composed of:
@@
RILe’s trainer continuously updates the reward function in tandem with the student’s policy updates, where IRL refines its reward function only after training a policy to convergence. Specifically, the trainer queries a discriminator to measure how expert-like the student’s behavior is, then optimizes the reward function based on that feedback, without waiting for the policy to converge. RILe preserves nuanced reward shaping, while avoiding IRL’s heavy computational loop. As a result, RILe is particularly effective in high-dimensional settings, where agents need fine-grained guidance at every stage of learning.
Our contributions are two-fold:
@@
    Efficient Reward-Function Learning via RL: We introduce a reinforcement-learning-based approach for training a reward function simultaneously with the policy. This avoids IRL’s repeated policy re-training and the purely discriminator-based rewards of AIL/AIRL. Reinforcement learning enables the trainer agent to explore multiple reward regimes, yielding a reward function that considers long-horizon effects.
@@
We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL, specifically GAIL [5] AIRL [8], GAIfO [9], BCO [10], IQ-Learn [7] and DRAIL [11]. Our experiments spans four scenarios: (1) Empirically investigating how RILe’s reward-learning strategy differs from AIL and AIRL in a maze task, (2) Quantifying the adaptability of the learned reward function in a humanoid locomotion task, (3) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems, and (4) Analyzing the impact of explicit expert-data usage within RILe’s training process. Our results show RILe’s superior performance, particularly in high-dimensional environments, and highlight RILe’s ability to learn a dynamic reward function that effectively guides the student through multiple stages of training.
@@ 2 Related Work
Early work in IL introduced Behavioral Cloning (BC) [4], which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations. DAgger [12] refines BC by aggregating data over multiple iterations to mitigate compounding errors. GAIL [5] employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior. BCO [10] extends BC, and GAIfO [9] extends GAIL, both to handle state-only observation scenarios. DQfD [13] introduces a two-stage approach with pre-training, while ValueDice [14] aligns policy and expert distributions via a distribution-matching objective. More recently, DRAIL [11] leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
Despite these advances, IL methods face challenges in high-dimensional environments [6, 7], where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance. RILe addresses these limitations through an adaptive trainer–student framework, where a learned reward function provides more nuanced guidance than standard IL comparison mechanisms.
@@
Inverse Reinforcement Learning (IRL), introduced by [1], aims to uncover the expert’s intrinsic reward function from demonstrations. Major developments include Apprenticeship Learning [15], Maximum Entropy IRL [2], and adversarial variants like AIRL [8]. IQ-Learn [7] reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability. More recent work focuses on unstructured data [16] and cross-embodiment transfer [17].
Nonetheless, IRL methods struggle with computational inefficiency and limited scalability [18], particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly. RILe mitigates these challenges by jointly learning the policy and reward function in a single process, avoiding IRL’s iterative retraining loop and facilitating more efficient reward shaping for complex environments.
@@ 3.1 Markov Decision Process
A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$. $S$ is the state space consisting of all possible environment states $s$, and $A$ is action space containing all possible environment actions $a$. $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function. $T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$. $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor. The policy $\pi=\pi(a|s):S\ \rightarrow\ A$ is a mapping from states to actions. In this work, we consider $\gamma$-discounted infinite horizon settings. Following [5], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$. The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$. We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [19].
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy, $\pi^{*}$. that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig. 1-(a)). In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ [5, 20]. The RL problem with a parameterized reward function and entropy regularization is defined as:
@@
Figure 1: Overview of the related works. (a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data. IRL has two stages: 1. training a policy with frozen reward function, and 2. updating the reward function by comparing the converged policy with data. These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function. GAIL trains both policy and the discriminator at the same time. AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green). (D) RILe: similar to IRL, learning a reward function from data. RILe learns the reward function at the same time with the policy, using a discriminator as a guide for learning the reward.
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig. 1-(b)). Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $\mathbb{E}_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})+H(\pi)]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}$. Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$. Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [2] is defined as
@@ 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
GAIL [5] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)). The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert ones $(s,a)\notin\tau_{E}$. The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint. The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
@@
AIRL [8] extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as:

### v4

@@ Abstract
Acquiring complex behaviors is essential for artificially intelligent agents, yet learning these behaviors in high-dimensional settings poses a significant challenge due to the vast search space. Traditional reinforcement learning (RL) requires extensive manual effort for reward function engineering. Inverse reinforcement learning (IRL) uncovers reward functions from expert demonstrations but relies on an iterative process that is often computationally expensive. Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons often offer insufficient feedback for effective learning. We introduce RILe (Reinforced Imitation Learning), a framework that combines the strengths of imitation learning and inverse reinforcement learning to learn a dense reward function efficiently and achieve strong performance in high-dimensional tasks. RILe employs a novel trainer–student framework: the trainer learns an adaptive reward function, and the student uses this reward signal to imitate expert behaviors. By dynamically adjusting its guidance as the student evolves, the trainer provides nuanced feedback across different phases of learning. Our framework produces high-performing policies in high-dimensional tasks where direct imitation fails to replicate complex behaviors. We validate RILe in challenging robotic locomotion tasks, demonstrating that it significantly outperforms existing methods and achieves near-expert performance across multiple settings.
@@ 1 Introduction
Learning complex behaviors is critical for advancing artificially intelligent agents in fields such as robotics and strategic games. Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design. This is both time-consuming and error-prone.
@@
There are two ways to address the reward engineering problem. First, Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000; Ziebart et al., 2008) offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering. IRL proceeds iteratively: it first trains a policy (the learning agent’s decision-making mechanism) using the current reward function, observes how well the agent’s behavior aligns with the expert’s, and then refines the reward function to better guide the policy toward expert-like behaviors. Repeating this process eventually yields a reward function capable of providing nuanced feedback at different stages of learning. However, this iterative procedure is computationally expensive (Zheng et al., 2022), especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
@@
Second, Imitation learning (IL) bypasses explicit reward design by directly comparing learned behaviors to expert demonstrations via a comparison mechanism. Traditional IL approaches such as Behavioral Cloning (BC) (Bain & Sammut, 1995) match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks. To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL (Ho & Ermon, 2016), introduce a discriminator as a comparison mechanism that judges how expert-like the learned behaviors are. However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements, or boosts exploration. Instead, they rely on a distance measure or a binary classifier, that merely checks whether the agent’s behavior is (or is not) similar to the expert. Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize, or which actions to explore. Consequently, both traditional IL and AIL struggle in high-dimensional environments (Peng et al., 2018; Garg et al., 2021), where the agent needs more granular and adaptive feedback than these mechanisms provide.
@@
Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2018) attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator. However, AIRL tightly couples the reward function to the discriminator’s output, causing it to inherit AIL’s limitations in high-dimensional settings where more fine-grained guidance is needed.
@@
In contrast, real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog. The teacher also refines how they teach as the student progresses. Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student more effectively. By contrast, existing approaches lack this cooperative synergy. Adversarial Imitation Learning (AIL) does update a discriminator alongside the policy, but the discriminator’s sole role is to distinguish expert-like behavior from non-expert behavior. Consequently, the student attempts to fool this judge into classifying its behavior as expert-like, resulting in a competitive process rather than a cooperative trainer that dynamically shapes rewards based on suboptimal behaviors. Meanwhile, IRL methods only refine the reward after the policy converges, missing the opportunity for continuous co-evolution throughout training.
@@
To address these issues, inspired by these insights, we propose Reinforced Imitation Learning (RILe). RIle combines the adaptive reward benefits of IRL with the computational efficiency of AIL (Fig. 1-(d)). RILe is a novel trainer-student system that establishes a positive-sum relationship between the trainer and the student. Specifically, RILe is composed of:
@@
RILe’s trainer continuously updates the reward function in tandem with the student’s policy updates, where IRL refines its reward function only after training a policy to convergence. Specifically, the trainer queries a discriminator to measure how expert-like the student’s behavior is, then optimizes the reward function based on that feedback, without waiting for the policy to converge. RILe offers nuanced reward shaping, while avoiding IRL’s heavy computational loop. As a result, RILe is particularly effective in high-dimensional settings, where agents need fine-grained guidance at every stage of learning. Our contributions are two-fold:
@@
    Efficient Reward-Function Learning via RL: We introduce a reinforcement-learning-based approach for training a reward function simultaneously with the policy. This avoids IRL’s repeated policy re-training and the purely discriminator-based rewards of AIL/AIRL. Unlike the competitive judge in AIL, using RL allows RILe’s trainer agent to explore reward strategies and learn adaptively from the student’s progress, establishing a cooperative dynamic that yields a reward function that considers long-horizon effects.
@@
We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL: GAIL (Ho & Ermon, 2016) AIRL (Fu et al., 2018), GAIfO (Torabi et al., 2018b), BCO (Torabi et al., 2018a), IQ-Learn (Garg et al., 2021) and DRAIL (Lai et al., 2024). Our experiments span six studies: (1) Empirically analyzing how RILe’s reward-learning differs from baselines, (2) Quantitatively analyzing the learned reward function in RILe, (3) Comparing different trainer-discriminator relationships in RILe, (4) Evaluating the noise robustness of RILe, (5) Analyzing the impact of using expert-data explicitly inside RILe, and (6) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems. Our results show RILe’s superior performance, particularly in high-dimensional environments, and highlight RILe’s ability to learn a dynamic reward function that effectively guides the student through multiple stages of training.
@@ 2 Related Work
Early work in IL introduced Behavioral Cloning (BC) (Bain & Sammut, 1995), which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations. DAgger (Ross et al., 2011) refines BC by aggregating data over multiple iterations to mitigate compounding errors. GAIL (Ho & Ermon, 2016) employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior. BCO (Torabi et al., 2018a) extends BC, and GAIfO (Torabi et al., 2018b) extends GAIL, both to handle state-only observation scenarios. DQfD (Hester et al., 2018) introduces a two-stage approach with pre-training, while ValueDice (Kostrikov et al., 2020) aligns policy and expert distributions via a distribution-matching objective. More recently, DRAIL (Lai et al., 2024) leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
Despite these advances, IL methods face challenges in high-dimensional environments (Peng et al., 2018; Garg et al., 2021), where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance. RILe addresses these limitations through an adaptive trainer–student framework, where a learned reward function provides more nuanced guidance than standard IL comparison mechanisms.
@@
Inverse Reinforcement Learning (IRL), introduced by Ng & Russell (2000), aims to uncover the expert’s intrinsic reward function from demonstrations. Major developments include Apprenticeship Learning (Abbeel & Ng, 2004), Maximum Entropy IRL (Ziebart et al., 2008), and adversarial variants like AIRL (Fu et al., 2018). IQ-Learn (Garg et al., 2021) reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability. More recent work focuses on unstructured data (Chen et al., 2021) and cross-embodiment transfer (Zakka et al., 2022).
Nonetheless, IRL methods struggle with computational inefficiency and limited scalability (Arora & Doshi, 2021), particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly. RILe mitigates these challenges by jointly learning the policy and reward function in a single process, avoiding IRL’s iterative retraining loop and facilitating more efficient reward shaping for complex environments.
@@ 3.1 Markov Decision Process
A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$. $S$ is the state space consisting of all possible environment states $s$, and $A$ is action space containing all possible environment actions $a$. $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function. $T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$. $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor. The policy $\pi=\pi(a|s):S\ \rightarrow\ A$ is a mapping from states to actions. In this work, we consider $\gamma$-discounted infinite horizon settings. Following Ho & Ermon (2016), expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$. The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$. We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ (Finn et al., 2016).
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks an optimal policy, $\pi^{*}$. that maximizes the discounted cumulative reward from the reward function $R=R(s,a)$ (Fig. 1-(a)). In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ (Ho & Ermon, 2016; Bloem & Bambos, 2014). The RL problem with a parameterized reward function and entropy regularization is defined as
@@
Figure 1: Overview of the related works. (a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data. IRL has two stages: 1. training a policy with frozen reward function, and 2. updating the reward function by comparing the converged policy with data. These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function. GAIL trains both policy and the discriminator at the same time. AIRL implements a new structure on the discriminator, separating reward from environment dynamics by using two networks under the discriminator (see additional terms in green). (D) RILe: similar to IRL, learning a reward function from data. RILe learns the reward function at the same time with the policy, using a discriminator as a guide for learning the reward.
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig. 1-(b)). Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $\mathbb{E}_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})+H(\pi)]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}$. Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$. Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning (Ziebart et al., 2008) is defined as
@@ 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
GAIL (Ho & Ermon, 2016) introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)). The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert ones $(s,a)\notin\tau_{E}$. The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint. The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
@@
AIRL (Fu et al., 2018) extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as: